### v1

# MalPurifier: Enhancing Android Malware Detection with Adversarial Purification against Evasion Attacks

Abstract
Machine learning (ML) has gained significant adoption in Android malware detection to address the escalating threats posed by the rapid proliferation of malware attacks. However, recent studies have revealed the inherent vulnerabilities of ML-based detection systems to evasion attacks. While efforts have been made to address this critical issue, many of the existing defensive methods encounter challenges such as lower effectiveness or reduced generalization capabilities. In this paper, we introduce a novel Android malware detection method, MalPurifier, which exploits adversarial purification to eliminate perturbations independently, resulting in attack mitigation in a light and flexible way. Specifically, MalPurifier employs a Denoising AutoEncoder (DAE)-based purification model to preprocess input samples, removing potential perturbations from them and then leading to correct classification. To enhance defense effectiveness, we propose a diversified adversarial perturbation mechanism that strengthens the purification model against different manipulations from various evasion attacks. We also incorporate randomized “protective noises” onto benign samples to prevent excessive purification. Furthermore, we customize a loss function for improving the DAE model, combining reconstruction loss and prediction loss, to enhance feature representation learning, resulting in accurate reconstruction and classification. Experimental results on two Android malware datasets demonstrate that MalPurifier outperforms the state-of-the-art defenses, and it significantly strengthens the vulnerable malware detector against 37 evasion attacks, achieving accuracies over 90.91%. Notably, MalPurifier demonstrates easy scalability to other detectors, offering flexibility and robustness in its implementation.
Index Terms:
Android Malware Detection, Machine Learning, Evasion Attacks, Adversarial Purification, Denoising Autoencoder.
1 Introduction
Due to the popularity of the Android operating system, it has become the primary victim of malware attacks. In 2021, Zimperium reported that 2 billion new malware emerged in the wild [1], and Kaspersky detected 1,661,743 mobile malware or unwanted software installers in 2022 [2]. As a result, the prevalence of Android malware has grown exponentially in recent years, posing a significant threat to the security and privacy of mobile users worldwide.
The magnitude of this threat has spurred the use of Machine Learning (ML) techniques, particularly Deep Learning (DL), to automate Android malware detection. Empirical evidence has shown that these approaches offer advanced performance in detecting malware (see, e.g., [3, 4, 5, 6, 7]), making them a promising avenue for mitigating this security concern.
Unfortunately, ML-based detectors are vulnerable to adversarial examples, which are created by modifying non-functional instructions in executable programs of existing malware [8, 9, 10]. Adversarial examples can enable a range of attacks, including evasion attacks [11, 12, 13], poisoning attacks [14, 15, 16], or a combination of both [17]. In this study, we specifically narrow our focus to evasion attacks, which are designed to deceive ML-based detection during the testing phase.
So far, adversarial training-based methods have shown great potential to safeguard ML models from evasion attacks [18, 19, 20]. By augmenting the training dataset with generated adversarial samples, adversarial training can increase the robustness of the trained model in future use. However, these methods still have certain disadvantages, including high computational costs [21] and a significant sacrifice in accuracy on clean data [22]. Furthermore, its effectiveness is strongly influenced by the similarity between the adversarial examples employed during the training and testing phases [23]. This may lead to overfitting of the model to specific perturbations, thereby negatively impacting its ability to generalize and detect unseen attacks.
Another defense technique, known as adversarial purification [24, 25, 26], aims to remove potential perturbations from input samples, resulting in purified samples that can be correctly classified by the target classifier. The purification model is usually trained independently of the classification model and does not necessarily require class labels [27]. As a result, it can mitigate unseen threats in a plug-and-play manner without re-training the target classifier [28], leading to less training overhead and more flexible employment. Nevertheless, existing purification solutions for image classification are not easily applicable to Android malware detection due to significant differences. The input space of the image classifier is continuous, while malware features are often discrete [29]. Hence, it is more challenging to develop a robust Android malware detector.
Inspired by the complementary strengths of existing approaches in image classification, we propose a customized adversarial purification-based method to defeat evasion attacks in Android malware detection, as illustrated in Fig. 1. This novel mechanism has the following highlighted features that enable it to eliminate perturbations in adversarial malware and restore it closer to its original form: (i) We first generate diversified adversarial malware samples by adding perturbations in a progressively increasing manner to represent varying evasion attacks. This approach enables the purification model to fully acquire the capability to effectively purify these perturbed samples. (ii) We further injected “protective noises” into the benign samples during the training of the purification model. This step enhances the distinction between benign and adversarial samples, preventing excessive purification of benign samples. (iii) Based on these generated samples with their original versions, we then train a Denoising AutoEncoder (DAE) based purification model with a customized loss function in an unsupervised manner, allowing us to accurately pre-process the samples for classification purposes. (iv) More importantly, the proposed method is non-intrusive, computationally efficient, and easily scalable to different detection models. The main contributions of this paper can be summarized as follows:
  • Attack mitigation with adversarial purification. We propose a robust Android malware detection method, termed MalPurifier, that leverages a plug-and-play DAE model to purify adversarial malware. To the best of our knowledge, this is the first attempt to use adversarial purification to mitigate evasion attacks in the Android ecosystem.
  • Trade-off between robustness and detection accuracy. Our proposed mechanism injects a diversified range of perturbations into malware samples, ranging from no injection to worst-case perturbations, to enhance robustness against different types of attacks. Moreover, we also implement safeguards to prevent false positives on benign samples during the purification process.
  • Accurate sample recovery via denoising autoencoder. We establish a DAE-based purification model for purifying adversarially perturbed samples, which is independent of the label space and the detection model. Especially, we incorporate reconstruction loss and prediction loss to help the model handle complex and noisy data, leading to better feature representation and sample recovery.
  • Experimental validation on public datasets. We compare MalPurifier with the state-of-the-art (SOTA) methods via Drebin [30] and Androzoo [31] datasets. Experimental results show that MalPurifier significantly outperforms other defenses against attacks with less overhead required. Finally, we release the source code at https://github.com/SEU-ProactiveSecurity-Group/MalPurifier.
The remainder of this paper is organized as follows. Section 2 reviews some preliminaries and Section 3 presents the problem formulation. Section 4 elaborates the methodology, and the experimental results are presented in Section 5. We explore the limitations of our work and open challenges in Section 6, and discuss related work in Section 7. Finally, conclusion and future work are summarized in Section 8.
Figure 1: Illustration of MalPurifier pre-processing samples via adversarial purification, before feeding them into the malware detector. This method projects the various perturbed samples back to their original forms, while reserving the feature representation of clean data, successfully striking a balance between robustness and accuracy without necessitating any changes to the architecture or parameters of the detection model.
2 Preliminaries
2.1 ML-based Malware Detection
The ML-based Android malware detection can be briefly described as follows. Formally, let $\mathcal{Z}$ be the problem space, and $z\in\mathcal{Z}$ be an Android application sample. In the context of machine learning, there will be a feature extraction function $\phi:\mathcal{Z}\rightarrow\mathcal{X}$ which maps the problem space into the feature space, where $\mathcal{X}\subset\mathbb{R}^{d}$ is a $d$-dimensional discrete space.
The Android malware detection can be usually treated as a binary classification, thus, let $f:\mathcal{Z}\rightarrow\mathcal{Y}$ be the malware detector that maps the problem space to the label space $\mathcal{Y}=\{0,1\}$, where ”0” (or ”1”) means that corresponding example is benign (or malicious), respectively. Additionally, let the malware detector use an ML model $\varphi_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta$ represents the model’s parameters. Therefore, we can conclude that $f(\cdot)=\varphi_{\theta}(\phi(\cdot))$.
Given a sample-label pair $(z,y)$ and ML-based malware detector $f$, we then have $x=\phi(z)$. We can easily achieve the prediction $f(z)$ and compare it with the ground-truth label $y$ to analyze the accuracy. To improve the detection accuracy, the main task of Android malware detection is to achieve the optimal parameters as follows.
$\theta^{*}\in\arg\min_{\theta}\mathbb{E}_{(z,y)\in\mathcal{D}}[\mathcal{L}(\theta,x,y)],$ (1)
where $\mathcal{L}(\theta,x,y)$ is the loss function for the ML model $\varphi_{\theta}$, and $\mathcal{D}$ represents the underlying data distribution of training examples.
2.2 Evasion Attacks
2.2.1 Attack Principle
According to Ref. [8], evasion attack can be categories into two types: problem-space attacks and feature-space attacks. In the problem-space attack, the adversary perturbs a malware sample from $z$ to $z^{\prime}$ to evade the detector $f$. Accordingly, they can be mapped into the feature space with $x=\phi(z)$ and $x^{\prime}=\phi(z^{\prime})$. Formally, given a feature-label pair $(x,y)$ of a malware sample and an adversarial manipulation $\delta$, the evasion attack can be written as
$\varphi_{\theta}(x^{\prime})=\varphi_{\theta}(x+\delta)=0,\quad\mathrm{s.t.}(x^{\prime}\in\mathcal{X})\land(x^{\prime}\in[\check{u},\hat{u}]),$ (2)
where $x^{\prime}$ is the perturbed feature representation. Recent studies have suggested that it obeys a box constraint [17], such that $x^{\prime}\in[\check{u},\hat{u}]$, where $\check{u}$ and $\hat{u}$ denote the lower and upper boundaries in the feature space, respectively.
In addition, to establish the inverse mapping from the feature space to the problem space, we adopt the methodology proposed in Ref. [32]. This approach facilitates the design of attack tactics while maintaining the effectiveness of the attack. By utilizing an approximate inverse function $\tilde{\phi}^{-1}$, we can directly map the perturbation vector $\delta$ to the problem space.
2.2.2 Attack Methods
The attack method defines how the attacker implements malicious actions. In this section, we consider four distinct attack methods as follows.
Obfuscation Attacks. This kind of attacks suggests malware authors leveraging obfuscation technology to camouflage malicious functionality [33, 11]. Typically, adversaries exploit certain techniques (e.g., encryption, renaming, etc.) to produce malware variants that can deceive detection. Note that this attack does not require knowledge of the target classifier, making it a zero-query black-box attack that can be directly performed on the problem space.
Gradient-based Attacks. These attacks apply small perturbations in the direction of gradients to produce adversarial malware samples. For example, Projected Gradient Descent (PGD) attack [34] initializes the perturbation with a zero vector and perturbs it via an iterative process, such that
$\delta^{t+1}=\Pi_{[\check{u}-x,\hat{u}-x]}\Big{(}\delta^{t}+\lambda\nabla_{\delta}\mathcal{L}(\theta,x+\delta^{t},y)\Big{)},$ (3)
where $t$ is the iteration, $\lambda>0$ is the step size, $\Pi_{[\check{u}-x,\hat{u}-x]}$ is the projection operator that keeps $\delta^{t+1}$ within a set of range $[\check{u}-x,\hat{u}-x]$, and $\nabla_{\delta}$ indicates the gradient of the loss function $\mathcal{L}$ with respect to $\delta$. Due to the small magnitudes of gradients in practical scenarios, researchers have been motivated to normalize the gradients in a direction of interest, such as the $\ell_{1}$, $\ell_{2}$, or $\ell_{\infty}$ norm [35].
Furthermore, this study incorporates several other algorithms to perform gradient-based attacks, including Bit Coordinate Ascent (BCA) [36], Fast Gradient Sign Method (FGSM) [19], and Grosse [37].
Gradient-free Attacks. These attacks are permitted to get access to a surrogate dataset and wage evasion attacks via perturbations. The salt and pepper noises attack [13] involves manipulating malware samples by randomly replacing feature values with either the maximum or minimum intensity values, resembling the spread of salt and pepper particles. This study also investigates the use of pointwise attacks [38], in which the adversary first adds noise perturbation and then modifies features to generate an adversarial sample with the least perturbation.
Ensemble Attacks. These attacks provide attackers with the capability to compromise the victim via a combination of multiple attack methods and manipulations. For instance, Li et al. [20] proposed a series of ensemble-based attacks, including the ”Max” strategy enabled Mixture of Attacks (MaxMA), iterative MaxMA (iMaxMA), and Stepwise Mixture of Attacks (StepwiseMA), which effectively enhance the attack performance. Additionally, Croce and Hein [39] combined powerful attacks to create an ensemble attack namely AutoAttack, which demonstrates strong generalization across different models.
2.3 Adversarial Purification
The fundamental concept behind adversarial purification is to preprocess the input data directly, preventing any embedded adversarial components from feeding into the target model, so that the influence of attacks can be mitigated. These methods are widely regarded as model-agnostic and highly efficient, making them easy to train and utilize while demonstrating strong generalization capabilities.
Let $g$ be the adversarial purifier that uses a generative model $\psi_{\vartheta}$ with $g(\cdot)=\psi_{\vartheta}(\phi(\cdot))$ to learn the data distribution closer to the training distribution and restore an adversarial example to its corresponding clean example, where $\vartheta$ represents its parameters. Given $x=\phi(z)$ and $x^{\prime}=\phi(z^{\prime})$, thus, the training objective of purification is then
$\vartheta^{*}\in\arg$ $\min_{\vartheta}\mathbb{E}_{(z,y)\in\mathcal{D}}[\mathcal{J}(\vartheta,x^{\prime},x)],$ (4)
$\mathrm{s.t.}(x^{\prime}\in\mathcal{X})$ $\land(x^{\prime}\in[\check{u},\hat{u}])\land(\psi_{\vartheta}(x^{\prime})\in\mathcal{X}),$
where $\mathcal{J}(\vartheta,x^{\prime},x)$ represents the loss function for the learning model $\psi_{\vartheta}$, $x$ denotes the original feature, and $x^{\prime}=x+\delta$ is the perturbed feature representation. This training procedure only focuses on the differences of the representation between the sample after purification and its original version, thus, we can clearly conclude that the purification model is trained independently of the class label.
Unlike adversarial example attacks in the image domain that perturb images with inconspicuous noises, adversaries in this field specifically employ discrete manipulations on malware samples to evade detection. These adversarial examples closely resemble benign data in their feature representation, posing a significant challenge for the purification model. For example, clean data might be mistaken for adversarial examples, resulting in incorrect restoration into samples with malicious feature representation. As a consequence, the accuracy of clean data may experience a substantial decrease. Thereby, it remains a question of how to effectively enhance the trade-off between robustness and accuracy of adversarial purification.
3 Problem Formulation
Here we first introduce the threats considered in our work, and then propose a defense formulation to guide the design of the robust Android malware detection method.
3.1 Threat Model
To perform analysis on potential attacks and develop countermeasures, we consider the threat model in terms of explicit assumptions regarding the attacker’s capabilities and knowledge of the target system. Especially, we discuss prevailing attack scenarios specified by three different threat models in this section.
3.1.1 Black-Box Attacks
The attacker has no knowledge about the malware detector $f$ or the purifier $g$, but can only obtain the prediction results from the target system. Given a malware example $z$, we can have $x=\phi(z)$, and the attacker aims to modify it for producing wrong prediction such that
$\varphi_{\theta}(\psi_{\vartheta}(x^{\prime}))=0,\quad\mathrm{s.t.}(x^{\prime}\in\mathcal{X})\land(x^{\prime}\in[\check{u},\hat{u}]),$ (5)
3.1.2 Grey-Box Attacks
The adversary has knowledge of the original malware classifier $f$ but is not aware of the purifier $g$ (i.e., oblivious attack). Hence, the adversary’s goal is only to fool the unsecured model $f$ using adversarial examples generated from malicious examples. Formally, given a malware example $z$ with its feature representation $x=\phi(z)$ and label $y=1$, the attacker will modify it to obtain the adversarial feature representation $x^{\prime}$ that can evade detection, by solving
$\max_{x^{\prime}\in[\check{u},\hat{u}]}\mathcal{L}(\theta,x^{\prime},1),\quad\mathrm{s.t.}x^{\prime}\in\mathcal{X},$ (6)
where we substitute $\varphi_{\theta}(x^{\prime})=0$ with maximizing $\mathcal{L}(\theta,x^{\prime},1)$ owing to the non-differentiability of $\phi(\cdot)$.
3.1.3 White-Box Attacks
In this attack scenario, the adversary has full knowledge of the model $f$ and is aware of the structure of the purifier $g$ (i.e., adaptive attack). In other words, the adversary needs to mislead both the malware detector and adversarial purifier simultaneously by perturbing $x$ into $x^{\prime}$. It is worth noting that the sample will be first feed into the purifier and its output will be then classified by the malware detector.
Thereby, given a malware feature-label pair $(x,y)$ in this study, the attacker needs to solve
$\max_{x^{\prime}\in[\check{u},\hat{u}]}\mathcal{L}(\theta,\psi_{\vartheta}(x^{\prime}),1),\quad\mathrm{s.t.}(x^{\prime}\in\mathcal{X})\land(\psi_{\vartheta}(x^{\prime})\in\mathcal{X}),$ (7)
where $\psi_{\vartheta}(x^{\prime})$ denotes the purified sample obtained by applying the adversarial purifier $g$ to the input $x^{\prime}$.
3.2 Defense Formulation
As aforementioned, MalPurifier is rooted in adversarial purification, aiming to eliminate potential adversarial manipulations before detection. Thereby, we propose incorporating the detector $f$ with the adversarial purifier $g(\cdot)=\psi_{\vartheta}(\phi(\cdot))$. To develop the framework of MalPurifier, we need to train the detection model $\varphi_{\theta}$ according to Eq. (1) and build the purification model $\psi_{\vartheta}$ based on Eq. (4), respectively. To this end, given a feature-label pair $(z,y)$ with possible adversarial perturbations $\delta$ in the feature space, the desired parameters $\theta^{*}$ and $\vartheta^{*}$ of MalPurifier can be derived by solving the following problem
$\varphi_{\theta^{*}}(\psi_{\vartheta^{*}}(x$ $+\delta))=y,$ (8)
$\mathrm{s.t.}(x\in\mathcal{X})\land(y\in$ $\mathcal{Y})\land(x+\delta\in[\check{u},\hat{u}]),$
when $\delta=0$, it indicates that no adversarial manipulations have been applied to this sample, rendering it clean. The above formulation points out two tasks as follows.
  • High-accuracy Android malware detection. Developing a malware detection model that can classify clean samples into benign or malicious with high accuracy, that is, $\varphi_{\theta}(x)=y$, in which $x$ can be the feature representation of a normal Android application or a malware sample, and thus $y=0$ or $1$. This model can be easily obtained by utilizing some ML algorithms (i.e., Deep Neural Network (DNN)), which have shown promising results with 99% accuracy in their laboratory settings [40, 12, 6].
  • Effective adversarial manipulation elimination. This formulation highlights the importance of effectively integrating adversarial purification with the pre-trained detector, as summarized into two aspects. (i) Given the feature representation of a malware sample $x$ with its adversarial version $x^{\prime}$, it is crucial to minimize the impact of evasion attacks by achieving $\psi_{\vartheta}(x^{\prime})\thickapprox x$. This enables us to accurately identify it through the malware detector, indicated by $\varphi_{\theta}(\psi_{\vartheta}(x^{\prime}))=1$. (ii) When dealing with a clean feature-label pair $(x,y)$, it is essential to preserve the accuracy on it. This entails ensuring that the prediction results of clean data remain unaffected, represented by $\varphi_{\theta}(\psi_{\vartheta}(x))=y$.
4 The MalPurifier Approach
In this section, we provide an in-depth exploration of the MalPurifier approach. We begin by offering an overview that encompasses the processes of model training and inference. Subsequently, we delve into the mechanisms of diversified adversarial perturbation and protective noise injection, which are employed to generate training data for the purification model. Finally, we elaborate on the establishment of a label-independent DAE-based purification model, which plays a crucial role in producing purified samples and thereby contributing to accurate predictions.
Figure 2: Overview of MalPurifier architecture. In the training phase, feature vectors are extracted from Android apps in Step 1. Then, a detection model is constructed using features from both benign and malicious apps in Step 2. In Step 3, diversified adversarial perturbations are applied to malware samples in the feature space, while projective noises are introduced into benign samples in Step 4. Finally, in Step 5, the purification model is built using these variant samples along with their corresponding original versions. In the testing phase, a sample undergoes sequential processing by the purifier and detector, ensuring that adversarial malware cannot escape detection.
4.1 Defense Architecture
The MalPurifier plays two important roles, i.e., adversarial purification and malware detection. First, it performs input preprocessing in a label-independent manner and generates reconstructed representations in new forms. Especially, adversarial samples are restored to their original versions, while the characteristics of clean data remain unchanged. Second, it determines the benignity of the inputs by analyzing the prediction results of the DNN model on the processed samples. Fig. 2 depicts an overview of the MalPurifier architecture and explains how the functionally different components are jointly optimized.
As we can see, the proposed method consists of three main modules: (i) a feature extractor $\phi(\cdot)$ that maps an Android application into a feature vector, (ii) an adversarial purifier $g$ that processes samples via a DAE model $\psi_{\vartheta}$, and (iii) a malware detector $f$ that uses a DNN model $\varphi_{\theta}$ for detection. In the training phase, we initiate the process by extracting features from multiple batches of clean (or natural) examples, which include both benign and malicious examples without manipulation. These features are then fed into the DNN model, which iteratively updates its parameters to minimize the loss function. Once the DNN model is effectively trained, it exhibits exceptional classification accuracy when presented with clean inputs. At the same time, as a detection model, it often requires a substantial amount of training samples. To reduce costs, we will not retrain this model in this study, in contrast to adversarial training methods.
Different from the DNN, the DAE learns compressed representations of input data and reconstructs the original (clean) data from noisy versions. The performance of a DAE model is significantly influenced by the training data, especially the level and type of noise introduced during training. To enhance the DAE’s ability to handle various adversarial perturbations in input data, we propose a diversified adversarial perturbation mechanism, as demonstrated in Section 4.2. This mechanism generates a range of adversarial malware samples with varying attack strengths, thereby improving the model’s effectiveness in mitigating evasion attacks. Furthermore, in this study, we introduce a noise injection mechanism (see Section 4.3) to safeguard benign samples from excessive or erroneous purification by the DAE, hence preserving their original feature patterns. Finally, we incorporate reconstruction loss and prediction loss into the DAE model, and feed both generated adversarial samples and noisy benign samples, along with their corresponding clean forms into it. Note that we train the DAE model independent of the label space. More details of the DAE training can be found in Section 4.4.
In the phase of testing, the MalPurifier can be fed with benign, malicious and adversarial examples. By inputting an Android App into our method, one can obtain the prediction result (benign or malicious) as it passes through these modules. As depicted in the right part of Fig. 2, the MalPurifier demonstrates accurate identification of adversarial samples that may have evaded detection without purification.
4.2 Diversified Adversarial Perturbation
To train an adversarial purification model, we need to create perturbed malware samples first. Traditional adversarial training methods usually exploit strong white-box attacks to consider the target parameters $\theta$ and perturb the malware sample to create $x^{\prime}$, such that it will be misclassified by the detection model, i.e., $\varphi_{\theta}(x^{\prime})=0$. Since this method calculates perturbations using gradient directions specific to $\theta$ and stops manipulations when the classifier achieves misclassification, these perturbed malware samples cannot generalize well across different attack intensities.
To solve the aforementioned challenges, we design a diversified adversarial perturbation mechanism that can generalize across multiple kinds of attacks, thus enabling a robust defense approach. Specifically, this mechanism aims at increasing the feature space difference between a malware sample $x$ and its perturbed version $x^{\prime}$ instead of specific label information, which can be represented as
$\Delta(x,x^{\prime})=d(\mathcal{F}_{\theta}(x)|_{n},\mathcal{F}_{\theta}(x^{\prime})|_{n}),\ \ \mathrm{s.t.}x^{\prime}\in[\check{u},\hat{u}],$ (9)
where $\mathcal{F}_{\theta}(x)|_{n}$ denotes the internal representation of input data $x$ from the $n$th layer of the ML-based malware detector $f$, and $d(\cdot)$ quantifies the difference between the features of the original and perturbed samples. In this study, we employ Mean Square Error (MSE) as the distance metric. Furthermore, instead of directly maximizing this gap under constrained conditions, we add perturbations in a gradually increasing manner during the processing of input samples. This approach allows the generated samples to encompass a wider range of attack intensities, ensuring diversified and highly generalizable adversarial examples.
Algorithm 1 demonstrates the generation of diversified adversarial examples as the following steps. (i) We first sample a malware batch $(x_{i},y_{i}=1)_{i=1}^{N}$ in Line 2. (ii) Then, Line 3 sets the adversarial depth to a linear relationship with the batch index. (iii) In line 5, we do not inject any perturbations for the first batch of malware samples. (iv) For the other batch, we randomly transform $x_{i}$ as the initial point in Line 7. We derive the gradients $g_{t}$ in Line 10 according to the difference calculated in Line 9, thereby creating adversarial samples according to the preset depth in Line 11. To keep the generated adversarial samples practical, we also project them into the binary space in Line 12. After $T$ rounds of iterations, we craft adversarial malware samples for this batch in Line 13. (v) We repeat the above steps on each batch of the dataset to obtain enough diversified adversarial examples as candidates for adversarial purification.
Input: Training dataset $(\mathcal{X},\mathcal{Y})$, number of batches $N$, number of iterations $T$, step size $s$, and random transformation function $\mathcal{R}$;
Output: Generated adversarial subset $\mathcal{X}_{adv}$;
1 for $i=1$ to $N$ do
2       Sample a batch of $(x_{i},y_{i}=1)$ from $(\mathcal{X},\mathcal{Y})$;
3       Adversarial depth $k_{i}=s*(i-1)$;
4       if $k_{i}=0$ then
5             $x_{i}^{\prime}\leftarrow x_{i}$; $\triangleright$ First batch without perturbation
6
7      else
8             $x_{i,1}^{\prime}=\mathcal{R}(x_{i})$; $\triangleright$ Create a random initial point
9             for $t=1$ to $T$ do
10                   Compute $\Delta$ between $x_{i}$ and $x_{i,t}^{\prime}$ via Eq.(9);
11                   Compute gradients $g_{t}=\nabla_{x_{i}}\Delta(x_{i},x_{i,t}^{\prime})$;
12                   Generate adversarial samples by $x_{i,t+1}^{\prime}=x_{i,t}^{\prime}+k_{i}*g_{t}$;
13                   Project $x_{i,t+1}^{\prime}$ into the binary space;
14
15            $x_{i}^{\prime}\leftarrow x_{i,T}^{\prime}$; $\triangleright$ Other batches with perturbation
16
17
Return $\mathcal{X}_{adv}=\{x_{1}^{\prime},x_{2}^{\prime},...,x_{N}^{\prime}\}$
Algorithm 1 Diversified Adversarial Perturbation
4.3 Protective Noise Injection
The proposed diversified adversarial perturbation mechanism successfully enhances the diversity of adversarial malware samples within the training dataset, however, if benign samples are left unprocessed, the purification model may have a tendency to misclassify them as malicious samples, resulting in erroneous classification outcomes. It is also important to note that adversaries typically do not alter benign samples to make them appear as malware. As a result, the direct application of adversarial perturbation methods to manipulate benign samples is not viable.
To solve this problem and protect clean samples from excessive purification, we here propose a projective noise injection mechanism that preprocesses clean data before the training phase of the purification model. Specifically, we introduce a threshold $\eta\in[0,1]$ to control the range of noise injection. When $\eta$ is set to 0, no noise is added, while a value of 1 indicates that all feature values will be flipped under this condition. By employing this approach, the purification model is encouraged to better discriminate between benign and malicious samples within the feature space, thereby ensuring the detection performance of subsequent classifiers on clean data.
The detailed algorithm of the protective noise injection is described in Algorithm 2. (i) For each batch of training dataset, we first select benign samples $(x_{i},y_{i}=0)$ in Line 2, where $i\in[1,N]$. (ii) Then, we calculate the batch size $b_{i}$ and length $l_{i}$ of this batch of benign data for the following noise injection. (iii) We randomly generate a mask and determine the injection position according to the preset noise level $\eta$. (iv) We modify binary feature values on these positions (flipping from ”0” to ”1” and from ”1” to ”0”) based on these generated masks. (v) Finally, by applying the aforementioned steps to all batches of the training dataset, we obtain a subset of benign samples that have been augmented with protective noise.
Input: Training dataset $(\mathcal{X},\mathcal{Y})$, number of batches $N$, number of iterations $T$, and noise level $\eta$;
Output: Processed benign subset $\mathcal{X}_{ben}$;
1 for $i=1$ to $N$ do
2       Sample a batch of $(x_{i},y_{i}=0)$ from $(\mathcal{X},\mathcal{Y})$;
3       Obtain its batch size $b_{i}$ and length $l_{i}$;
4       Generate random mask $m=rand(b_{i},l_{i})<\eta$;
5       Flip feature values via $x_{i}[m]=1-x_{i}[m]$;
6
Return $\mathcal{X}_{ben}=\{x_{1},x_{2},...,x_{N}\}$
Algorithm 2 Protective Noise Injection
4.4 Accurate Sample Recovery
In this section, we introduce the technical details of the proposed purification mechanism and explain how to effectively recover perturbed adversarial malware samples while not affecting the detection of clean data through this mechanism. Our purification model is built on a denoising autoencoder, which has excellent capability to remove noise and improve the quality of the data for downstream tasks.
As illustrated in Fig. 2, the input of the purification model $\psi_{\vartheta}$ is the perturbed sample generated from diversified adversarial perturbation in Algorithm 1 (see Step 3), the modified benign samples created from protective noise injection in Algorithm 2 (see Step 4), as well as their original forms. During the training process of the purification model, these data are passed through and the network parameters $\vartheta$ can be optimized to minimize the customized loss that is described below.
Reconstruction Loss. A commonly used loss function for DAEs is the MSE loss, which measures the average squared difference between two groups of data. To help train a more robust purification model, we aim to minimize the discrepancy of the feature representation between the reconstructed output and the target data (i.e., reconstruction loss) as follows.
$\mathcal{L}_{rec}=d(x,\psi_{\vartheta}(x^{\prime})),\ \ \mathrm{s.t.}x^{\prime}\in\mathcal{X}_{adv}\cup\mathcal{X}_{ben},$ (10)
where $x$ and $\psi_{\vartheta}(x^{\prime})$ denote the original data and the purified data, respectively.
Prediction Loss. Given that the ultimate objective of adversarial purification is to enhance the classification performance of the subsequent malware detector, we propose using prediction loss to further optimize the performance of the purification model as follows.
$\mathcal{L}_{pre}=\Delta(x,\psi_{\vartheta}(x^{\prime}))=d(\mathcal{F}_{\theta}(x)|_{n},\mathcal{F}_{\theta}(\psi_{\vartheta}(x^{\prime}))|_{n}),$ (11)
where $\Delta$ is formally defined in Eq. (9), and the distance measure used to compute $\Delta$ is MSE as well. In this way, the purification model will be strengthened in which the purified data will be closer to the original version in the internal representation of the prediction model (i.e., the malware detector), leading to more accurate prediction results.
Therefore, the overall loss function for $\psi_{\vartheta}$ is the combination of reconstruction loss and prediction loss as follows.
$\mathcal{L}_{\psi_{\vartheta}}=\alpha\mathcal{L}_{rec}+\beta\mathcal{L}_{pre},\ \ \mathrm{s.t.}\alpha,\beta\in[0,1],$ (12)
where $\alpha$ and $\beta$ are the weights of two loss terms, and we have $\alpha+\beta=1$. The parameters $\vartheta$ will be optimized by minimizing the customized loss of the purification model.
5 Experiments and Evaluation
In this section, we conduct extensive experiments to validate the soundness of MalPurifier by answering the following Research Questions (RQs):
  • RQ1: Effectiveness and cost without attacks. How is the effectiveness and overhead of MalPurifier when there is no attack?
  • RQ2: Robustness against obfuscation attacks. How is the robustness of MalPurifier against obfuscation attacks?
  • RQ3: Robustness against oblivious attacks. How robust is MalPurifier against oblivious attacks where the attacker is unaware of the additional defense mechanism (e.g., the adversarial purifier $g$)?
  • RQ4: Robustness against adaptive attacks. How robust is MalPurifier against adaptive attacks in which the adversary has full knowledge of all defense mechanisms?
  • RQ5: Enhancement brought to other detectors. Can the purification model act as a plug to enhance other detectors against evasion attacks?
Datasets. Our experiments utilize two popular Android malware datasets: Drebin [30] and Androzoo [31]. The Drebin dataset [^1] https://www.sec.cs.tu-bs.de/ danarp/drebin consists of 5,560 malicious samples and SHA256 values of 123,453 benign applications, which were collected before 2013. For evaluation purposes, we downloaded 47,770 benign APKs from various markets (e.g., Google Play Store, AppChina, Anzhi). To obtain more recent files, we collected 170,851 APKs from the AndroZoo dataset [^2] https://androzoo.uni.lu, specifically those attached with dates falling between January 1st and December 31st, 2021. We submitted these APKs to the VirusTotal [^3] https://www.virustotal.com/gui/home/upload service, labeling a sample as malicious if at least five anti-virus scanners raised alarms, and considering it benign if no scanner detected it. We randomly selected 10,987 benign examples and 10,998 malicious examples from Androzoo for our experiments. Note that each dataset was randomly split into three distinct sets for training (60%), validation (20%), and testing (20%).
Feature extraction. Drebin [30] analyzes a set of APKs and constructs a suitable feature space. Thus, we here utilize the Androguard [^4] https://github.com/androguard/androguard tool to perform a static analysis and extract the Drebin features, which can be organized in 8 different feature sets, including 4 subsets extracted from the manifest (e.g., hardware components), and the other 4 subsets extracted from the disassembled dexcode (e.g., API calls). The APK is mapped into the feature space as a binary feature vector, in which we can have 0 or 1 along each dimension, indicating the presence or absence of the corresponding feature. Following prior work [20], we exclude certain features that can be easily renamed or modified (e.g., package name) and retain the most frequent 10,000 ones in this study.
Defenses considered for comparative analysis. In this paper, we compare the SOTA defense mechanisms as follows:
  • DNN [41]. It employs a DNN model for malware detection without any countermeasures against evasion attacks.
  • ${}^{+}$ [42]. It enhances the robustness of the detector by another detector trained with an additional outlier class for detecting adversarial examples.
  • KDE [43]. It introduces a secondary detector that utilizes a Kernel Density Estimate (KDE) method. This detector identifies adversarial examples in the final layer of the DNN that deviate significantly from normal data.
  • FD-VAE [23]. It improves the DNN model by introducing an additional Variational AutoEncoder (VAE) for Feature Disentangle (FD) in different classes and combining their detection outcomes to make the final decision (FD-VAE).
[…]
  • AT-Adam [44]. It enhances the robustness of the DNN model via incorporating Adversarial Training with the PGD attack optimized by Adam (AT-Adam).
  • PAD-SMA [20]. It achieves Principled Adversarial Detection by a DNN-based malware detector and an Input Convexity Neural Network (ICNN) based adversary detector, both of which are strengthened by adversarial training incorporating the Stepwise Mixture of Attacks (PAD-SMA).
All above defenses consider DNN as the baseline classifier, they either improve the malware detector via adversarial training, or introduce another detector to identify adversarial examples, or both of them. Unlike these methods, MalPurifier takes a different approach. Firstly, it avoids the use of adversarial training methods to retrain the malicious detector, opting instead to fix the detector to minimize costs. Secondly, to guide the detector towards accurate classification, MalPurifier focuses on removing potential perturbations from the samples rather than attempting to detect adversarial samples as a new class.
Metrics. The effectiveness of defenses is assessed using five standard metrics: False Positive Rate (FPR), False-Negative Rate (FNR), Accuracy (Acc), balanced Accuracy (bAcc), and F1 score. The inclusion of balanced Accuracy and F1 score accounts for the imbalanced nature of the dataset. In addition, we include training time to evaluate the overhead of these methods.
5.1 RQ1: Effectiveness and Cost without Attacks
Experimental Setup. We compare MalPurifier with the aforementioned approaches on the two datasets. We use a DNN model with 2 fully-connected hidden layers (each having 200 neurons) with the ELU activation, and the other methods also use this architecture for malware detection.
[…]
The proposed method exploits a DAE-based purification model, which has two layers (each having 600 neurons) for both the encoder and decoder with the Sigmoid activation and introduces attention weights following the encoder. The training data of the DAE model is generated by Algorithm 1 with step size 0.01 and Algorithm 2 with noise level 0.001. We conduct a group of preliminary experiments and finally set $\alpha=\beta=0.5$ on both datasets. In addition, all detectors are tuned by the Adam optimizer with 100 epochs, batch size 128, and learning rate 0.001.
TABLE I: Effectiveness and overhead of detectors in the absence of attacks.
Defense Effectiveness (%) Overhead (s)
FPR FNR Acc bAcc F1 Training time
Drebin DNN 0.51 8.07 98.72 95.71 93.58 592
${}^{+}$ 0.54 7.79 98.72 95.83 93.60 1559
KDE 0.53 8.10 98.67 95.68 93.53 592
FD-VAE 1.14 23.3 96.62 87.79 82.12 1021
[…]
AT-Adam 4.01 5.47 95.84 95.26 82.14 1341
PAD-SMA 1.70 5.94 97.87 96.18 89.93 21627
MalPurifier 2.55 7.05 97.00 95.20 86.23 750
Androzoo DNN 0.32 1.22 99.23 99.23 99.23 527
${}^{+}$ 0.05 0.83 99.54 99.56 99.56 1148
KDE 0.15 1.22 99.31 99.32 99.32 527
FD-VAE 11.26 4.37 92.22 92.19 92.55 2182
[…]
AT-Adam 3.54 0.59 97.95 97.94 98.00 1018
PAD-SMA 0.90 1.62 98.73 98.74 98.77 29837
MalPurifier 2.30 0.59 98.57 98.56 98.59 696
Results. Table I exhibits the effectiveness and overhead of all detectors when there is no attack. We observe that ${}^{+}$ achieves the highest detection accuracy (98.72% on Drebin and 99.23% on Androzoo) and F1 score (93.58% on Drebin and 99.56% on Androzoo), which are a little higher than those of the basic DNN model. The reason may be that adversarial training introduces extra adversarial examples that help ${}^{+}$ identify more malicious samples, resulting in a lower FNR and higher F1 score.
[…]
Answer to RQ1: MalPurifier exhibits a slight decrease in accuracy on clean data with a slightly increased overhead. In comparison, FD-VAE experiences a significant decrease in accuracy whereas PAD-SMA has an excessively long training time on both datasets.
Figure 3: The accuracy of different detectors against obfuscation attacks on Drebin and Androzoo datasets. The color gradient ranging from light to dark represents the increasing accuracy from low to high, with the effectiveness of MalPurifier against each attack annotated in the square.
5.2 RQ2: Robustness against Obfuscation Attacks
Experimental Setup. We measure the accuracy of all aforementioned methods under obfuscation attacks via the Drebin and Androzoo datasets. Since these attacks belong to black-box attacks and can be conducted in the problem space, therefore, we leverage obfuscation technology to modify malicious APKs and extract features from the modified versions. Specifically, we utilize an obfuscator called AVPASS [33], to wage 8 obfuscation-based attacks to perturb malware examples on the test set.
For Java Reflection (JF), this attack can hide public and static system APIs invoked in Smali using the reflection API. The encryption attacks typically encrypt the const-string and variable names in the decode, i.e., String Encryption (SE) and Variable Encryption (VE). The Package Name Change (PNC), Method Name Change (MNC), and Class Name Change (CNC) attacks change the names of packages, methods, and classes by replacing them with random characters, respectively. For the Resource Obfuscation (RO) attack, it changes pixel or adds one byte to the image files of APKs, along with the modification of related AndroidManifest.xml. Finally, we combine the above techniques to produce a Combined Attack (CA).
Results. Fig. 3 illustrates the accuracy of the detectors on Drebin (left panel) and Androzoo (right panel) datasets under 8 obfuscation attacks. We make the first observation that DNN can not defeat all these attacks (accuracy $\leqslant$ 0.344% on Drebin and $\leqslant$ 5.871% on Androzoo), demonstrating that such attacks can hide malicious features to evade detection. Nevertheless, obfuscation attackers produce adversarial examples in a black-box manner, they cannot effectively evade these detectors except for the DNN model.
[…]
Answer to RQ2: MalPurifier outperforms the other methods against all obfuscation attacks on the Drebin dataset (accuracy of 100%), and achieves accuracy $\geqslant$ 95% against 7 obfuscation attacks on the Androzoo dataset.
5.3 RQ3: Robustness against Oblivious Attacks
Figure 4: The accuracy of different detectors against gradient-based oblivious attacks on Drebin (top panel) and Androzoo (bottom panel) datasets, along with the iteration ranging from 0 to 100.
[…]
We separately wage 12 oblivious attacks on test malware examples. For BCA [36], Grosse [37], and PGD-$\ell_{1}$ [35] attacks, we perturb one feature per time with a maximum iteration of 100. For rFGSM [19] and $\ell_{\infty}$ [35], we iterate these attack algorithms with the iterations 100 and step size 0.02. The PGD-$\ell_{2}$ [35] attack is set with 100 iterations and step size 0.5. Moreover, we perform Salt & Pepper attack [13] by increasing the noise intensity of 0.001 each time until misclassification and repeating this process 10 times, and the Pointwise [38] attack utilizes Salt & Pepper as the initial attack and minimizes the needed perturbations.
Additionally, we combine PGD-$\ell_{1}$, PGD-$\ell_{2}$, and $\ell_{\infty}$ attacks to perform MaxMA [20] attack and run it 5 times with the random starting point for the iMaxMA [20] attack. We also iterate the SMA [20] attack 100 times with the step size 0.5 for PGD-$\ell_{2}$ and 0.02 for $\ell_{\infty}$, and the AutoAttack [39] consists of APGD-CE and FAB with $\ell_{2}$ norm in our experiments.
Results. Fig. 4 depicts the accuracy curves of these methods on Drebin (top panel) and Androzoo (bottom panel) datasets under 6 gradient-based oblivious attacks with the iteration from 0 to 100. We first observe an important observation that none of these attacks can evade MalPurifier (accuracy $\leqslant$ 90.91% on Drebin and $\leqslant$ 99.41% on Androzoo), demonstrating the high robustness of the proposed approach.
We further observe that there is a decreasing trend on the curves of ${}^{+}$, KDE, and PAD-SMA against BCA, Grosse, and PGD-$\ell_{1}$ attacks until 20 iterations. This is because such attacks in an oblivious manner will stop manipulations when the adversarial example can evade the malware detector. Moreover, there exists a dip in some accuracy curves of ${}^{+}$, KDE, and FD-VAE against rFGSM and PGD-$\ell_{2}$ attacks with the iteration ranging from 0 to 100. This is because such attacks create small perturbations that can successfully escape on a specific iteration, and as the number of iterations increases, they will be recognized due to the larger perturbations.
An interesting observation is that KDE can mitigate $\ell_{\infty}$ on both datasets whereas fails to defeat the other attacks. The underlying reason for this observation may be that KDE relies on the close distance between activations to reject large manipulations that are used by the $\ell_{\infty}$ attack, while the basic DNN model is very sensitive to small perturbations, leading to the failure against the other attacks. Another interesting observation is that the robust accuracies of FD-VAE against all these gradient-based attacks are very different on the two datasets. The reason may be that the threshold of reconstruction error in FD-VAE relies on the distribution of training datasets, which may be significantly different from each other.
TABLE II: Accuracy of different defenses against oblivious attacks in which attackers only know the original malware detector.
Attack Name Accuracy (%)
${}^{+}$ KDE FD-VAE PAD-SMA MalPurifier
Drebin No Attack 92.21 91.93 73.19 94.06 95.08
Salt & Pepper 0.000 0.000 100.0 100.0 100.0
Pointwise 0.000 0.000 69.46 89.70 100.0
MaxMA 24.58 91.93 58.63 94.25 96.66
iMaxMA 24.58 91.93 58.63 94.25 96.66
StepwiseMA 12.71 0.649 13.82 89.05 96.66
AutoAttack 81.35 82.93 42.30 93.69 96.94
Androzoo No Attack 98.74 98.83 95.59 98.20 99.41
Salt & Pepper 87.98 89.46 91.90 99.96 99.24
Pointwise 77.36 71.56 90.68 97.97 99.24
MaxMA 19.71 98.83 94.78 98.42 99.41
iMaxMA 19.71 98.83 94.78 98.42 99.41
StepwiseMA 61.21 8.236 18.23 77.32 99.41
AutoAttack 1.800 27.41 38.25 98.42 99.41
Table II reports the results of Salt &\&& Pepper, Pointwise, MaxMA, iMaxMA, StepwiseMA, and Autoattack, which are not suitable for a large number of iterations. We first observe that MalPurifier can effectively mitigate these attacks and achieves the highest accuracy against them, except for the Salt &\&& Pepper attack with 99.24% accuracy on Androzoo. The results indicate that our method significantly improves the robustness and outperforms the state-of-the-art methods in terms of gradient-free, gradient-based, and ensemble-based attacks.
We further observe that there exist significant differences in the accuracy values of all defensive methods when dealing with various attacks, except for MalPurifier. For example, KDE can effectively mitigate MaxMA and iMaxMA with an accuracy of 91.93% but cannot defeat the StepwiseMA attack with an accuracy of only 0.649% on Drebin. This is because the other methods are vulnerable to large or small manipulations, while the diversified adversarial perturbation mechanism in MalPurifier can help defend against a range of perturbations from different attacks. Note that the purification model is not dependent on any specific attacks, so all attacks in our experiments are unknown to MalPurifier, highlighting the advantages of this method.
Answer to RQ3: MalPurifier is significantly more robust than ${}^{+}$, KDE, FD-VAE, and PAD-SMA. It achieves an accuracy of $\geqslant$ 90.91% and $\geqslant$ 99.24% on both datasets against oblivious attacks, wherein all of these adversarial examples are previously unseen by the model.
TABLE III: Accuracy of different defenses under adaptive attacks where adversaries know all defensive mechanisms (if applicable).
Attack Name Accuracy (%)
[…]
Drebin BCA 0.000 6.122 42.21 0.000 0.000 5.102 80.15 90.91
rFGSM 0.000 14.94 85.81 59.46 91.93 74.40 94.25 95.08
Grosse 0.000 6.122 41.93 0.000 0.000 2.690 78.76 90.91
PGD-$\ell_{1}$ 0.000 0.186 41.84 0.000 0.000 0.835 77.83 99.72
PGD-$\ell_{2}$ 9.833 30.61 86.74 0.093 62.06 2.319 92.30 95.08
$\ell_{\infty}$ 0.000 15.21 92.30 51.67 91.93 75.97 94.25 95.08
Salt & Pepper 0.000 97.96 99.07 0.000 0.000 80.43 95.55 100.0
Pointwise 0.000 94.43 94.53 0.000 0.000 67.97 87.65 99.72
MaxMA 0.000 0.371 45.83 0.000 0.000 1.299 77.37 95.08
iMaxMA 0.000 0.371 45.83 0.000 0.000 1.299 77.83 95.08
StepwiseMA 0.000 0.371 45.55 0.371 25.05 1.206 88.22 95.08
AutoAttack 0.000 78.39 71.71 81.35 82.93 42.30 93.69 96.94
Orth PGD-$\ell_{1}$- -- -- 0.000 6.401 12.43 94.25 95.08
Orth PGD-$\ell_{2}$- -- -- 3.711 0.000 17.90 94.25 95.08
Orth $\ell_{\infty}$- -- -- 46.01 89.98 55.29 94.25 95.08
Orth MaxMA -- -- -- 0.000 0.000 13.08 94.25 95.08
Orth iMaxMA -- -- -- 0.000 0.000 13.08 94.25 95.08
Androzoo BCA 0.000 0.090 35.01 0.000 10.62 30.56 86.72 99.37
rFGSM 0.000 8.776 98.16 6.841 98.83 96.00 98.43 99.41
Grosse 0.000 0.090 35.01 0.000 0.045 28.04 56.30 99.37
PGD-$\ell_{1}$ 0.000 0.000 15.26 0.000 0.585 3.465 40.41 99.41
PGD-$\ell_{2}$ 62.96 89.96 93.70 89.83 73.58 90.14 98.34 97.75
$\ell_{\infty}$ 0.000 90.19 93.34 18.59 98.83 97.71 98.42 99.41
Salt & Pepper 89.47 99.73 99.82 87.17 89.06 91.85 98.38 99.41
Pointwise 71.56 99.25 99.33 76.50 68.36 90.68 94.27 99.37
MaxMA 0.000 0.000 14.94 0.000 0.585 7.111 40.41 97.44
iMaxMA 0.000 0.000 14.94 0.000 0.585 6.391 40.41 97.44
StepwiseMA 0.000 0.000 16.11 0.000 79.16 3.555 76.73 99.41
AutoAttack 0.540 98.25 97.80 1.800 27.41 38.25 98.42 99.41
Orth PGD-$\ell_{1}$- -- -- 0.000 0.000 8.731 98.42 99.41
Orth PGD-$\ell_{2}$- -- -- 0.045 78.22 43.70 98.42 99.41
Orth $\ell_{\infty}$- -- -- 19.13 93.65 97.21 98.42 99.41
Orth MaxMA -- -- -- 0.000 0.000 7.111 98.42 99.41
Orth iMaxMA -- -- -- 0.000 0.000 7.111 98.42 99.41
5.4 RQ4: Robustness against Adaptive Attacks
[…]
Results. Table III reports the experimental results against adaptive attacks. First, we can observe that DNN is very vulnerable to all attacks, with 0% accuracy against 11 attacks on Drebin and 8% accuracy against 8 attacks on Androzoo. However, the Salt &\&& Pepper and Pointwise attacks achieve the lowest attack effectiveness in evading DNN on Androzoo, because both of them modify malware examples without using the internal information of target detectors.
[…]
Third, although ${}^{+}$, KDE, and FD-VAE incorporate another adversary detector, they only show limited effectiveness under a few attacks, and suffer from unseen attacks such as PGD-$\ell_{1}$, MaxMA, iMaxMA, Orth PGD-$\ell_{1}$, Orth MaxMA, and Orth iMaxMA (with an accuracy $\leqslant$ 13.08% on both datasets). Additionally, PAD-SMA not only hardens the DNN with adversarial training, but also combines it with an ICNN-based adversary detector, significantly improving its robustness against different attacks. Especially, PAD-SMA achieves the highest accuracy of 98.34% against the PGD-$\ell_{2}$ attack on Androzoo. However, PAD-SMA is still sensitive to the Grosse, PGD-$\ell_{1}$, MaxMA, and iMaxMA attacks (with an accuracy $\leqslant$ 78.76% on Drebin and $\leqslant$ 56.30% on Androzoo). Considering its high time overhead reported in Table I, it cannot be called a perfect solution.
In summary, MalPurifier significantly outperforms other defenses, achieving the highest accuracy against all 17 adaptive attacks on the Drebin dataset and 15 attacks on the Androzoo dataset. This indicates that the purification model can accurately recover the original forms of adversarial examples even if the adversary knows its existence.
Answer to RQ4: MalPurifier outperforms other defenses in the condition that adversaries know all defensive mechanisms, and significantly hardens the malware detector against a wide range of adaptive attacks (with an accuracy $\geqslant$ 90.91% on Drebin and $\geqslant$ 97.44% on Androzoo).
5.5 RQ5: Enhancement Brought to Other Detectors
Figure 5: The accuracy of different detectors against various evasion attacks when equipped with (w/) or without (w/o) the DAE-based purification model on Drebin (drawn in red) and Androzoo (drawn in blue) datasets.
Experimental Setup. To evaluate the flexibility and generalization of the purification model, we also package the DAE model as a plug to protect other detectors, such as Support Vector Machine (SVM), Fully Convolutional Network (FCN), Long Short Term Memory (LSTM), and Recurrent Neural Network (RNN). In detail, We use an SVM model with the Sigmoid activation and an FCN model with 3 fully-connected hidden layers (each having 512, 256, and 128 neurons) with the ReLU activation. The LSTM model has a hidden layer with 200 neurons, a sequence length of 1, and uses the Sigmoid function as the activation. Furthermore, we build the RNN model with 3 hidden layers (each having 200 neurons) and the Sigmoid activation. All these models are tuned by the Adam optimizer with 100 epochs, batch size 128, and learning rate 0.001.
Note that the DAE-based purification model works as a plug-and-play preprocessing method, in which we do not retrain the model but directly apply it from the aforementioned experiments. In addition, we wage 6 attacks on test malware examples in an adaptive manner, including BCA, PGD-$\ell_{1}$, PGD-$\ell_{2}$, $\ell_{\infty}$, Pointwise, and StepwiseMA, with the same hyper-parameters in Section 5.4.
Results. Fig. 5 depicts the accuracy improvement of these methods equipped with the DAE-based purification model on Drebin (drawn in red) and Androzoo (drawn in blue) datasets under 6 adaptive attacks. We make three observations as follows.
First, all detectors without enhancement cannot mitigate these adaptive attacks (with accuracy $\leqslant$ 16.42% on Drebin and $\leqslant$ 8.19% on Androzoo), except for the PGD-$\ell_{2}$ attack. This is because ML-based detectors themselves are very vulnerable to these evasion attacks with small perturbations. Especially, the attack effectiveness of the Pointwise attack is 100% when adversaries wage attacks on original detectors. Nevertheless, the PGD-$\ell_{2}$ attack, if running with a large number of iterations, will produce larger perturbations that may not evade detection.
Second, the DAE-based purification model works very well and can significantly improve the robustness of these detectors as a security plug (accuracy increase by $\geqslant$ 39.61% for SVM, $\geqslant$ 87.71% for FCN, $\geqslant$ 30.47% for LSTM, and $\geqslant$ 7.97% for RNN). Significantly, it boosts the accuracy of the RNN model against the Pointwise attack from 0% to 99.72% on Drebin and from 0% to 99.82% on Androzoo, rendering the previously vulnerable RNN model robust against this specific attack. These results strongly demonstrate the versatility of the DAE model, as it can seamlessly transfer to other models without the need for retraining.
Third, the accuracy values of these detectors are similar against different evasion attacks when equipped with the DAE model. The underlying reason for the observation is that the DAE model is trained in an independent and unsupervised way, and can accurately return the adversarial examples to their original forms. Hence, the effectiveness of equipping this security plug relies more on the detector itself, as the DAE model solely preprocesses the input data while the detector is responsible for classification.
Answer to RQ5: The DAE model (the core module of MalPurifier) can be flexibly transferred to other detectors without retraining. Notably, it greatly enhances the robustness against evasion attacks, even when these models are highly vulnerable to such attacks.
6 Limitations and Open Issues
While MalPurifier has yielded highly promising results, it is evident that this approach possesses some limitations. We additionally conduct experiments to assess the robustness of MalPurifier against the Mimicry attack, wherein the attacker introduces perturbations to a malware sample to closely resemble a benign application. Unfortunately, MalPurifier cannot effectively resist Mimicry ($\leqslant$ 60%) and has less effectiveness as malware samples are guided by more benign samples. This is primarily because the DAE model tends to purify these samples as benign due to their similarity to benign samples, leading to their misclassification. To mitigate this issue, we believe that implementing countermeasures, such as generating a more diverse set of adversarial examples or enhancing the adversarial purification model itself, would be beneficial. We plan to explore these possibilities in our future research.
Another limitation of our approach may be the potential performance degradation as new malware samples evolve. Even within this study, the effectiveness on the Androzoo dataset differs from that on the Drebin dataset, partly due to the inclusion of new samples. This vulnerability does not lie within the MalPurifier framework itself. We believe that by incorporating new data and leveraging dynamic updates, we can enhance the detection system’s capabilities.
While our approach may not be foolproof, we firmly believe that it substantially enhances the resistance of Android malware detection against diverse evasion attacks in a lightweight and plug-and-play manner.
7 Related Work
This section begins with a review of existing studies on ML-based Android malware detection methods, followed by an introduction to evasion attacks against these approaches and a brief discussion of state-of-the-art solutions.
7.1 ML-based Android Malware Detection
Researchers have developed numerous ML-based Android malware detection methods that typically classify APKs using features extracted from the manifest and bytecode. For instance, Drebin[30] identifies Android malware by exploiting binary static features and employing SVM for classification. MaMaDroid[46] extracts sequences of API calls and then trains classifiers like K-Nearest Neighbors (KNN) to detect malware.
Additionally, DL-based methods [40, 47] have demonstrated remarkable capabilities. For example, Andre [48] is a hybrid representation learning approach that clusters Android malware from multiple sources and classifies them using a three-layer DNN when they behave like existing families. Qiu et. al [5] proposed a framework that extracts heterogeneous features and utilizes DNN to recognize unknown and evolving malware.
Given the widespread use and outstanding performance of DNNs in Android malware detection, this study aims to enhance the robustness of DNN-based detectors using an independently trained purifier to pre-process input samples. This purifier restores the feature representation of adversarial malware to its original version and preserves the features of clean samples as much as possible, enabling the DNN model to correctly classify Android applications.
7.2 Evasion Attacks in Android Malware Detection
In the context of Android malware detection, evasion attacks employ crafted inputs to mislead models such that malicious apps will be classified as benign. As discussed in Section 2, it can be divided into problem-space attacks and feature-space attacks.
Problem-space attacks modify the Android apps directly, such as perturbations onto Android manifest and Dalvik bytecode [12] or insertion of benign components into malicious samples [8], for generating adversarial malware to deceive ML-based detectors. On the contrary, feature-space attacks map the malware example into a feature vector, and then introduce perturbations to the vector values [29] or reconstruct the vector representation [41] to achieve misclassification. Moreover, recent studies demonstrate that the utilization of ensemble attacks [49, 13] intensifies the impact of the attacks, presenting a more formidable challenge for defense mechanisms.
To combat the escalating prevalence of evasion attacks, the method presented in this paper is not tailored to counter any specific attack. Instead, it strives to establish a universally applicable approach that effectively mitigates both problem-space attacks and feature-space attacks. Additionally, the proposed method significantly enhances robustness while maintaining accuracy on clean samples.
7.3 Defenses against Evasion Attacks
Adversarial training [18, 22, 50] is widely recognized as one of the most popular methods for defeating evasion attacks. Recent research [13, 51] has further shown that combining adversarial training with ensemble learning can enhance model robustness. However, it is worth noting that adversarial training typically retrains the model by generating and incorporating adversarial examples, which can lead to a significant increase in computational burden. Also, the defenses may not effectively mitigate attacks that differ significantly from the ones encountered during training.
In addition, there are also several countermeasures to identify evasion attacks through an auxiliary model. For example, Li et. al [23] introduced a Variational AutoEncoder (VAE) to distinguish benign examples from adversarial malware according to reconstruction errors, and Li et. al [20] leverages a convex DNN model-based detector to recognize the evasion attacks. Despite not requiring retraining of the target model, the auxiliary model remains closely coupled with the malware detection model and is still unable to effectively handle sophisticated and adaptive attacks.
Different from existing studies, the proposed method involves diversified perturbations to increase its generalization capabilities against novel or unseen attacks. Furthermore, the independent training of the purification model makes it computationally efficient and easily scalable.
8 Conclusion and Future Work
We devise a plug-and-play defense method against evasion attacks for malware detection. Instead of hardening the malware detector solely, an independently training purification model is used to purify adversarial malware. In the training phase, we generate adversarial examples via injecting diversified perturbations into malware, along with protective noises added onto benign samples, to achieve a trade-off between robustness and detection accuracy. Additionally, we establish a DAE-based purification model with a customized loss function, resulting in better feature representation and robust accuracy. Experiments via two Android datasets demonstrate the soundness of the method against a set of attacks, including unseen ones.
Given the current trend in the use of diffusion models for adversarial purification, the future development of our work, which may further improve classifier security, is to leverage this technology to defend against evolving evasion attacks. Another interesting future extension of our approach may be to investigate robust methods against poisoning attacks, including the purification samples in the both training and test phases. These two parts of the research will substantially improve the security of employing machine learning techniques in Android malware detection.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62202097 and 62072100, in part by the China Postdoctoral Science Foundation under Grant 2022M710677, and in part by the Jiangsu Funding Program for Excellent under Grant 2022ZB137.